40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13252     Entry time: Fri Aug 25 01:20:52 2017     In reply to: 13248     Reply to this: 13276
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locking attempt 

I tried some DRMI locking again tonight, but had no success. Here is the story.

  • I started out by going to the AS table and measuring the light level on the REFL55 photodiode (with PRM aligned and the PRC flashing, but LSC disabled).
    • The Ophir power meter reads 13mW
    • The DC output of the photodiode shows ~500mV on an oscilloscope.
    • Both of these numbers line up well with measurements I made in April/May.
  • Returned to the control room and aligned the IFO for DRMI locking - but LSC servos remained disabled.
    • At the nominal REFL55 whitening level of +18dB, the REFL 55 signals saturated the ADC (confirmed by looking at the traces on dataviewer).
    • But the signals still looked like PDH error signals.
    • Lowering the whitening gain to 6dB makes the PDH error signal horns peak around 20,000 counts.
    • Could this be indicative of problems with either the analog whitening gain switching or the LSC Demod Boards? To be investigated.
  • Tried enabling LSC servos with same settings with which I had success right up till a couple of months ago, but had no success.
    • If it is true that the REFL55 signal is getting amplified because of some gain stage not being switched correctly, I should still have been able to lock the SRC with a lowered loop gain - but even lowering the gain by a factor of 10 had no effect on the locking success rate.

Looks like I will have to embark on the REFL55 LSC electronics investigation. I was able to successfully lock the PRC on carrier and sideband, and the Michelson lock also seems to work fine, all of which seem to point to a hardware problem with the REFL55 signal chain.

I did a quick check by switching the output of the REFL55 demod board to the inputs normally used by AS55 signals on the whitening board. Setting the whitening gain to +18dB for these channels had the same effect - ADC overflow galore. So looks like the whitening board isn't to blame. I will have to check the demod board out.

 

ELOG V3.1.3-